OpenStudy (anonymous):

23. Is the relationship between the variables in the table a direct variation, an inverse variation, or neither? If it is a direct or inverse variation, write a function to model it. x 8 11 14 17 y 1 10 18 31

OpenStudy (anonymous):

I got inverse varation; y=8/x am i right?

OpenStudy (ash2326):

@Math4Life \(y=\frac{8}{x}\) is satisfying only the first one. You see when x=11 y=10 and as x increases y is also increasing so it's not an inverse relation. Let's check if it's a direct relation

OpenStudy (anonymous):

direct varitation, y=1/8x ?

OpenStudy (ash2326):

Let's have the relation y=mx+c We'll use first two data and try to find m and c And then we'll check if it satisfies the others

OpenStudy (anonymous):

if it's not " direct variation, y=1/8x then it has to be neither

OpenStudy (ash2326):

From first we have x=8, y=1 so \[1=8m+c........................................(1)\] from second we have, x=11, y=10 \[10=11m+c..................................... (2)\] Let's subtract 1 from 2, we get \[10-1=11m+c-(8m+c)\] We get \[9=3m\] so \[m=3\] now we need to find c \[1=8m+c\] m=3 \[1=24+c\] so \[c=-23 \] so the relation is \[y=3x-23\] now let's check if it satisfies the other points

OpenStudy (anonymous):

ok please continue, thank you this is a really big help!

OpenStudy (ash2326):

Third x=14 and y=18. Let's check if it satisfies Our proposed relation \[y=3x-23\] x=14 \[y=3\times 14-23= 18\] Wow this one satisfied this. I've a feeling that this is our relation:D:D But we have to check others also

OpenStudy (ash2326):

Fourth x=17 y=31 \[y=3x-23\] x=17 \[y=3 \times 17-23\] \[y=51-23=28 \ne 31\] oh no:(:( This relation failed for this Unfortunately there is no relation between x and y. It's neither:(

OpenStudy (anonymous):

@ash2326 thank you so much! with showing your work i really understood it! thanks! now this is my first question so what do i do after you answer? do i give u a grade or points or something?

OpenStudy (ash2326):

You thanked me. That's enough. Welcome to Open Study:D I'm glad to help you:D:D

OpenStudy (amistre64):

if you dont see a "good answer" button next to ashes name; hit f5 to refresh your browser and see if it appears. the thank you is sufficient tho; but the medals are that little extra :)

OpenStudy (ash2326):

@Math4Life One important thing. Here at Open Study We have a code of conduct. Always follow it and you'd be a nice user. Here is the link to that http://openstudy.com/code-of-conduct Read it through Thanks

OpenStudy (ash2326):

@amistre64 I think this user left. Could you please close this question?

OpenStudy (anonymous):

@ash2326 Was I not being a nice user?

OpenStudy (ash2326):

You are a nice user. I think you won't have any problem following the code of conduct.:D

OpenStudy (anonymous):

@ash2326 Ok thanks! Well i'm going to close the question, thank you!

OpenStudy (ash2326):

Great:D:D